OpenStudy (anonymous):

Find the common ratio for the following numbers, then use the ratio to find the next three terms. 1) 1, -4, 16, -64 2) 9, 3, 1, 1/3 3) -3, 15, -75, 375 4) 24, 36, 54, 81

OpenStudy (jdoe0001):

in a geometric sequence, the common ratio or common multiplier, is just a factor, thus to find it you could always just divide the next term by the previous term, that is \(\bf 1st,\quad 2nd, \quad 3rd, \quad 4th\qquad \textit{common ratio}\implies \cfrac{2nd}{1st}\ or\ \cfrac{3rd}{2nd}\ or \cfrac{4th}{3rd}\)

OpenStudy (ranga):

I will do the first one and you can use the same procedure to answer the rest. 1) 1, -4, 16, -64 Common ratio = any term / previous term common ratio r = -4/1 = -4 or 16/-4 = -4 or -64/16 = -4 The next three terms are: -64 * -4 = 256; 256 * -4 = -1024; -1024 * -4 = 4096

OpenStudy (anonymous):

Thank you ranga that helps me a lot I really appreciate the help